OpenStudy (anonymous):

Find the radius of a cylinder with a volume of 188.5 feet³ and height of 15 feet

OpenStudy (anonymous):

\[V = pi r ^{2}h\] 188.5 = pi r^{2}15 pi r^2 = 188.5/15 = 12.57 r^2 = 12.57/pi = 4.000 r = 2